# Characterizing the support of semiclassical measures for higher-dimensional cat maps

### v1

@@ Appendix A Examining Orbit Closures
In this appendix, we characterize $K$ using the characteristic polynomial of $A$. Recall that $V_{v}\subset\mathbb{Q}^{2n}$ is the smallest rational subspace such that $\mathbb{R}v\subset V_{v}\otimes\mathbb{R}$ and that $\mathbb{T}_{v}\subset\mathbb{T}^{2n}$ is the subtorus given by the projection of $V_{v}\otimes\mathbb{R}$ onto $\mathbb{T}^{2n}$. We first characterize $\mathbb{T}_{v}$, adapting the argument from [dyatlov2021semiclassical]*Lemma 4.3.
@@
Set $G\coloneqq\overline{\mathbb{R}v\mod\mathbb{Z}^{2n}}$. As $\mathbb{R}v$ mod $\mathbb{Z}^{2n}\subset\mathbb{T}_{v}$, we know $G\subset\mathbb{T}_{v}$. It remains to show the reciprocal inclusion.
Since the set $G$ is a closed subgroup of $\mathbb{T}^{2n}$, it is a Lie subgroup. Let $\mathfrak{g}\subset\mathbb{R}^{2n}$ be the Lie algebra of $G$. Since $G$ is connected and abelian, the exponential map $\mathfrak{g}\rightarrow G$ is onto. Thus, $G\simeq\mathfrak{g}/Z$, where $Z\coloneqq\mathfrak{g}\cap\mathbb{Z}^{2n}$. As $G$ is compact, the rank of the lattice $Z$ is equal to the dimension of $\mathfrak{g}$. Therefore, letting $V\subset\mathbb{Q}^{2n}$ be the subspace generated by $Z$, $\mathfrak{g}=V\otimes\mathbb{R}$. As $v\in\mathfrak{g}$, by the definition of $V_{v}$, we have $V_{v}\subset V$. Thus, $V_{v}\otimes\mathbb{R}\subset\mathfrak{g}$, which implies $\mathbb{T}_{v}\subset G$. ∎
@@
$k_{0}\coloneqq\min\{k\in\mathbb{N}:\text{ the characteristic polynomial of }A^{k}\text{ is reducible over }\mathbb{Q}\}.$
@@
Set $Z_{i}\coloneqq X_{i}\oplus[0]_{l-m}$. As $[Z_{i},Z_{j}]=0$ and $\exp(sR)=\begin{bmatrix}\cos s&\sin s\\ -\sin s&\cos s\end{bmatrix}$, we see that $H_{0}\simeq\{e^{s_{1}Z_{1}+\cdots+s_{m}Z_{m}}:s_{1},\ldots,s_{m}\in\mathbb{R}\}$, where the isomorphism is given by a linear map. Thus, we see that $H_{0}v$ is isomorphic to a torus of dimension at most $m$.
3. Let $W$ be a rational hyperplane in $E_{+}$. We examine the intersection $W\cap H_{0}v$. As $H_{0}v$ is isomorphic via a linear transformation to a torus, either $H_{0}v\subset W$ or $W\cap H_{0}v$ has measure 0 in $H_{0}v$.
First suppose $H_{0}v\subset W$. By the definition of $H$, there exists some $L\geq 1$ for which $B^{L}\in H_{0}$. We assume that $L$ is the smallest such natural number for which this holds. Then, $A^{kL}v\in W$ for all $k\in\mathbb{Z}$. Let $W^{\prime}$ be the minimal rational subspace such that $A^{kL}v\subset W^{\prime}$ for all $k\in\mathbb{Z}$. Clearly, $W^{\prime}$ is $A^{L}$-invariant. As $W^{\prime}$ is a nontrivial proper rational subspace of $\mathbb{R}^{2n}$, $A^{L}$ must have reducible characteristic polynomial over $\mathbb{Q}$. Thus, $k_{0}\leq L$.
Therefore if $k_{0}=\infty$, then for all rational hyperplanes $W$ in $E_{+}$, $H_{0}v\cap W$ has measure 0 in $H_{0}v$. As the set of all rational hyperplanes in $E_{+}$ is countable, there exists $w\in H_{0}v$ that is not contained in any rational hyperplane. For some subsequence $k_{l}$, we have $B^{k_{l}}v\rightarrow w$. Passing to a further subsequence, there exists $z_{\infty}\in\mathbb{T}^{2n}$ such that $A^{k_{l}}z\bmod\mathbb{Z}^{2n}\rightarrow z_{\infty}$. Thus,
$\mathbb{T}^{2n}=\overline{\{z_{\infty}+\mathbb{R}w\bmod\mathbb{Z}^{2n}\}}\subset\overline{\{A^{k_{l}}(z+\mathbb{R}v)\bmod\mathbb{Z}^{2n}:l\in\mathbb{Z}\}}\subset K.$
Now suppose $k_{0}<\infty$. If $V_{v}=\mathbb{Q}^{2n}$, then $K=\mathbb{T}^{2n}$ and trivially $K$ contains the union of $k_{0}$ non-parallel tori of dimension at least 1. Otherwise, $11\leq\dim V_{v}\otimes\mathbb{R}\leq 2n-$ 1. Thus, for $0<k<k_{0}$, $V_{v}\otimes\mathbb{R}$ cannot be $A^{k}$-invariant. We then know for $0\leq k<k_{0}$, each torus $A^{k}\mathbb{T}_{v}$ has a different tangent space. We conclude
@@
i.e., $K$ contains the union of $k_{0}$ non-parallel tori of dimension at least 1. ∎
@@
$A$ is a hyperbolic element of $\operatorname{Sp}(4,\mathbb{Z})$ with $k_{0}=2$ and has an associated semiclassical measure supported exactly on the union of two tori.
@@
Let $\gamma^{\pm 1}$ be the eigenvalues of $B$, where $\gamma>1$. Therefore the eigenvalues of $A^{2}=-(B^{2}\oplus B^{2})$ are $-\gamma^{\pm 2}$. Thus, $i\gamma$, $-i\gamma$, $i\gamma^{-1}$, and $-i\gamma^{-1}$ are the eigenvalues of $A$, which implies that $A$ is hyperbolic.
@@
However, as the characteristic polynomial of $A^{2}=-(B^{2}\oplus B^{2})$ is the square of the characteristic polynomial of $-B^{2}$, it is clearly reducible over $\mathbb{Q}$. Therefore, $k_{0}=2$.
@@
Clearly, the characteristic polynomial of the above example in Proposition A.3 cannot have Galois group $S_{2}\wr S_{n}$. In fact, its Galois group is the Klein four group $V_{4}\simeq Z_{2}\times Z_{2}$. As noted earlier, the eigenvalues of $A$ are $i\gamma$, $-i\gamma$, $i\gamma^{-1}$, and $-i\gamma^{-1}$. Each of the maps $i\gamma\mapsto-i\gamma$, $i\gamma\mapsto i\gamma^{-1}$, and $i\gamma\mapsto-i\gamma^{-1}$ is an involution. Moreover, each of these maps swaps the other two roots; respectively, $i\gamma^{-1}$ swaps with $-i\gamma^{-1}$, $-i\gamma$ swaps with $-i\gamma^{-1}$, and $-i\gamma$ swaps with $i\gamma^{-1}$. Thus, every nontrivial element of the Galois group is of order two and a product of two disjoint transpositions. This is exactly the definition of the Klein four group as a subgroup of the symmetric group $S_{4}$.
@@ Appendix B Galois groups of random symplectic matrices
For any fixed natural number $n\geq 1$, $100\%$ of matrices $A\in\mathrm{Sp}(2n,\mathbb{Z})$ (when ordered by the maximum absolute value of their entries, say, or by any other norm on the space of $2n\times 2n$ matrices) are such that the characteristic polynomial of $A^{m}$ is irreducible with Galois group $S_{2}\wr S_{n}$ for every integer $m\geq 1$.
@@
Let $A\in\mathrm{Sp}(2n,\mathbb{Z})$ be such that the characteristic polynomial of $A$ is irreducible with Galois group $S_{2}\wr S_{n}$. If $n\geq 2$, then for every integer $m\geq 1$, the characteristic polynomial of $A^{m}$ is irreducible with Galois group $S_{2}\wr S_{n}$. If $n=1$, this same conclusion holds as well, provided that the characteristic polynomial of $A$ is not $x^{2}+1$, $x^{2}+x+1$, or $x^{2}-x+1$.
@@
Suppose to the contrary that $\lambda^{m}=\alpha$ for some $\alpha\in F$. The extension $\mathbb{Q}(\lambda)/F$ is Galois (it is degree $2$), with the nontrivial automorphism sending $\lambda$ to $\frac{1}{\lambda}$. Applying this automorphism (which fixes $\alpha$) and simplifying, we also find that $\lambda^{m}=1/\alpha$. This implies that $\alpha=\pm 1$ and that $\lambda$ is a root of unity. This cannot be the case if $n\geq 2$, since the extension $\mathbb{Q}(\lambda)$ is not abelian (the group $S_{2}\wr S_{n}$ is not abelian for any $n\geq 2$), nor can it be the case if $n=1$ unless $f_{A}$ is one of the listed polynomials (the listed polynomials are the only irreducible quadratics whose roots are roots of unity). This completes the proof. ∎
To prove Theorem B.1 in the case $m=1$, we will employ a strategy of Davis, Duke, and Sun [DDS], incorporating results of Chavdarov [Chavdarov] and Gorodnik and Nevo [GorodnikNevo]. We begin with the following lemma from [DDS].
@@
[DDS]*Lemma 2 Let $n\geq 2$ and let $G$ be a subgroup of $S_{2}\wr S_{n}$ containing a $2$-cycle, $4$-cycle, $(2n-2)$-cycle, and a $2n$-cycle. Then $G=S_{2}\wr S_{n}$.
Lemma B.3 suggests that sieving is a natural approach to Theorem B.1. Specifically, the characteristic polynomial $f_{A}(x)$ of any element $A\in\mathrm{Sp}(2n,\mathbb{Z})$ must be reciprocal (i.e., palindromic), that is, it satisfies $f_{A}(x)=x^{2n}f_{A}(\frac{1}{x})$; let $\mathcal{P}_{2n}^{\mathrm{rec}}(\mathbb{Z})$ be the set of monic reciprocal polynomials in $\mathbb{Z}[x]$ of degree $2n$, so that $f_{A}\in\mathcal{P}_{2n}^{\mathrm{rec}}(\mathbb{Z})$. Moreover, the factorization type of $f_{A}\,(\mathrm{mod}\,\ell)$ for primes $\ell$ not dividing the discriminant of $f_{A}$ corresponds to the cycle type of an element in $\mathrm{Gal}(f_{A})$ (namely, the Frobenius element at $\ell$). Thus, for any prime $\ell$, let $\mathcal{P}_{2n}^{\mathrm{rec}}(\mathbb{F}_{\ell})$ be the set of monic reciprocal polynomials of degree $2n$ over $\mathbb{F}_{\ell}$, and, motivated by Lemma B.3, for any $1\leq k\leq n$ we let $\mathcal{P}_{2n}^{\mathrm{rec}}(F_{\ell};(2k))$ be the subset of $\mathcal{P}_{2n}^{\mathrm{rec}}(\mathbb{F}_{\ell})$ consisting of polynomials factoring as an irreducible polynomial of degree $2k$ times a product of $2n-2k$ distinct linear polynomials. Observe that #$\#\mathcal{P}_{2n}^{\mathrm{rec}}(\mathbb{F}_{\ell})=\ell^{n}$.
@@
#$\#\mathcal{P}_{2n}^{\mathrm{rec}}(\mathbb{F}_{\ell};(2k))=\frac{1}{2^{n-k+1}\cdot k\cdot(n-k)!}\ell^{n}+O(\ell^{n-1}).$
@@
In particular, Lemma B.5 shows that the characteristic polynomials of elements of $\mathrm{Sp}(2n,\mathbb{F}_{\ell})$ are approximately equidistributed in $\mathcal{P}^{\mathrm{rec}}_{2n}(\mathbb{F}_{\ell})$. (Recall that $|\mathrm{Sp}(2n,\mathbb{F}_{\ell})|=\ell^{2n^{2}+n}+\linebreak O(\ell^{2n^{2}+n-1})$ and #$\#\mathcal{P}_{2n}^{\mathrm{rec}}(\mathbb{F}_{\ell})=\ell^{n}$.) However, to obtain Theorem B.1, it is necessary to know that elements of $\mathrm{Sp}(2n,\mathbb{Z})$ are essentially equidistributed when reduced $\,(\mathrm{mod}\,\ell)$, and that these reductions are essentially independent for different primes $\ell$. In other words, we are asking for an effective form of approximation for the group $\mathrm{Sp}(2n,\mathbb{Z})$. For this, we have the following consequence of work of Gorodnik and Nevo [GorodnikNevo].
@@
#$\#\{A\in\mathrm{Sp}(2n,\mathbb{Z}):\|A\|\leq H,A\equiv A_{0}\,(\mathrm{mod}\,q)\}$
@@
$\mathcal{E}_{2k}:=\{f\in\mathcal{P}_{2n}^{\mathrm{rec}}(\mathbb{Z}):f\,(\mathrm{mod}\,\ell)\not\in\mathcal{P}_{2n}^{\mathrm{rec}}(\mathbb{F}_{\ell};(2k))\text{ for all primes $\ell$}\}.$
@@
(B.1) #$\#\{A\in\mathrm{Sp}(2n,\mathbb{Z}):\|A\|\leq H,f_{A}\in\mathcal{E}_{2k}\}=o(\#\{A\in\mathrm{Sp}(2n,\mathbb{Z}):\|A\|\leq H\}).$
@@
#$\#\{A\in\mathrm{Sp}(2n,\mathbb{Z}):\|A\|\leq H,f_{A}\in\mathcal{E}_{2k}(q)\}$
@@
Acknowledgements. EK is supported by MathWorks and by NSF GRFP under grant No. 1745302. She would like to thank Semyon Dyatlov, who is partially supported by NSF CAREER grants DMS-1749858 and DMS-2400090, for suggesting and advising this project. She would also like to thank Bjorn Poonen and Vijay Srinivasan for their help with Proposition A.3. She is grateful to RJLO for his help with the Galois group explanation after Proposition A.3. TCA is supported by NSF DMS-2231990 and NSF CAREER DMS-2237937. She thanks Peter Sarnak for introducing a variant of this problem to her. RJLO is supported by NSF DMS-2200760 and a Simons Fellowship in Mathematics.

### v2

@@ Appendix A Examining Orbit Closures
In this appendix, we characterize $K$ using the characteristic polynomial of $A$. Recall that $V_{v}\subset\mathbb{Q}^{2n}$ is the smallest rational subspace such that $\mathbb{R}v\subset V_{v}\otimes\mathbb{R}$ and that $\mathbb{T}_{v}\subset\mathbb{T}^{2n}$ is the subtorus given by the projection of $V_{v}\otimes\mathbb{R}$ onto $\mathbb{T}^{2n}$. We first characterize $\mathbb{T}_{v}$.
@@
The argument is exactly the same as  [dyatlov2021semiclassical]*Lemma 4.3 and needs no adaptation to higher dimensions. ∎
@@
[…]
@@
Set $Z_{i}\coloneqq X_{i}\oplus[0]_{l-m}$. As $Z_{i},Z_{j}]=0$ and $\exp(sR)=\begin{bmatrix}\cos s&\sin s\\ -\sin s&\cos s\end{bmatrix}$, we see that $H_{0}\simeq\{e^{s_{1}Z_{1}+\cdots+s_{m}Z_{m}}:s_{1},\ldots,s_{m}\in\mathbb{R}\}$, where the isomorphism is given by a linear map. Thus, we see that $H_{0}v$ is isomorphic to a torus of dimension at most $m$.
3. Let $W$ be a rational hyperplane in $E_{+}$. Specifically, we mean that $W$ is given by a basis of l1l-1 elements of $\mathbb{Q}^{2n}\cap E_{+}$. We examine the intersection $W\cap H_{0}v$. As $H_{0}v$ is isomorphic via a linear transformation to a torus, either $H_{0}v\subset W$ or $W\cap H_{0}v$ has measure 0 in $H_{0}v$.
4. First suppose $H_{0}v\subset W$. By the definition of $H$, there exists some $L\geq 1$ for which $B^{L}\in H_{0}$. We assume that $L$ is the smallest such natural number for which this holds. Then, $A^{kL}v\in W$ for all $k\in\mathbb{Z}$. Let $W^{\prime}$ be the minimal rational subspace such that $A^{kL}v\subset W^{\prime}$ for all $k\in\mathbb{Z}$. Clearly, $W^{\prime}$ is $A^{L}$-invariant. As $W^{\prime}$ is a nontrivial proper rational subspace of $\mathbb{R}^{2n}$, $A^{L}$ must have reducible characteristic polynomial over $\mathbb{Q}$. Thus, $k_{0}\leq L$.
If $V_{v}=\mathbb{Q}^{2n}$, then $K=\mathbb{T}^{2n}$ and trivially $K$ contains the union of $k_{0}$ non-parallel tori of dimension at least 1.
Now assume that $V_{v}\neq\mathbb{Q}^{2n}$. In this case, $1\leq\dim V_{v}\otimes\mathbb{R}\leq 2n-1$. By the definition of $k_{0}$, for $0<k<k_{0}$, $A^{k}$ has no proper, nontrivial rational invariant subspaces. Therefore, $V_{v}\otimes\mathbb{R}$ cannot be $A^{k}$-invariant. In other words, $A^{k}(V_{v}\otimes\mathbb{R})\neq A^{j}(V_{v}\otimes\mathbb{R})$ for $0\leq k<j<k_{0}$. Then, for $0\leq k<k_{0}$, each torus $A^{k}\mathbb{T}_{v}$ has a different tangent space. We conclude
@@
i.e., $K$ contains the union of $k_{0}$ non-parallel tori of dimension at least 1.
5. Now suppose that $W\cap H_{0}v$ has measure 0 in $H_{0}v$. From the above argument, we see $k_{0}=\infty$. As the set of all rational hyperplanes in $E_{+}$ is countable, there exists $w\in H_{0}v$ that is not contained in any rational hyperplane. For some subsequence $k_{l}$, we have $B^{k_{l}}v\rightarrow w$. Passing to a further subsequence, there exists $z_{\infty}\in\mathbb{T}^{2n}$ such that $A^{k_{l}}z\bmod\mathbb{Z}^{2n}\rightarrow z_{\infty}$. Thus,
$\mathbb{T}^{2n}=\overline{\{z_{\infty}+\mathbb{R}w\bmod\mathbb{Z}^{2n}\}}\subset\overline{\{A^{k_{l}}(z+\mathbb{R}v)\bmod\mathbb{Z}^{2n}:l\in\mathbb{Z}\}}\subset K,$
which finishes the proof. ∎
@@
AA is a hyperbolic element of $\operatorname{Sp}(4,\mathbb{Z})$ with $k_{0}=2$ and has an associated semiclassical measure supported exactly on the union of two tori.
@@
Let $\gamma^{\pm 1}$ be the eigenvalues of $B$, where $\gamma>1$. Therefore the eigenvalues of $A^{2}=-(B^{2}\oplus B^{2})$ are $\gamma^{\pm 2}$. Thus, $i\gamma$, $i\gamma$, $i\gamma^{-1}$, and $i\gamma^{-1}$ are the eigenvalues of $A$, which implies that $A$ is hyperbolic.
@@
However, as the characteristic polynomial of $A^{2}=-(B^{2}\oplus B^{2})$ is the square of the characteristic polynomial of $B^{2}$, it is clearly reducible over $\mathbb{Q}$. Therefore, $k_{0}=2$.
@@
Clearly, the characteristic polynomial of the above example in Proposition A.3 cannot have Galois group $S_{2}\wr S_{n}$. In fact, its Galois group is the Klein four group $V_{4}\simeq Z_{2}\times Z_{2}$. As noted earlier, the eigenvalues of $A$ are $i\gamma$, $i\gamma$, $i\gamma^{-1}$, and $i\gamma^{-1}$. Each of the maps $i\gamma\mapsto-i\gamma$, $i\gamma\mapsto i\gamma^{-1}$, and $i\gamma\mapsto-i\gamma^{-1}$ is an involution. Moreover, each of these maps swaps the other two roots; respectively, $i\gamma^{-1}$ swaps with $i\gamma^{-1}$, $i\gamma$ swaps with $i\gamma^{-1}$, and $i\gamma$ swaps with $i\gamma^{-1}$. Thus, every nontrivial element of the Galois group is of order two and a product of two disjoint transpositions. This is exactly the definition of the Klein four group as a subgroup of the symmetric group $S_{4}$.
@@ Appendix B Galois groups of random symplectic matrices
For any fixed natural number $n\geq 1$, there exists a subset $\mathcal{S}$ of $\mathrm{Sp}(2n,\mathbb{Z})$ such that
$\lim_{H\to\infty}\frac{\#\{A\in\mathcal{S}:\|A\|\leq H\}}{\#\{A\in\mathrm{Sp}(2n,\mathbb{Z}):\|A\|\leq H\}}=1,$
every $A\in\mathcal{S}$ has a non-unit length eigenvalue, and so that for every $A\in\mathcal{S}$ and every integer $m\geq 1$, the characteristic polynomial of $A^{m}$ is irreducible with Galois group $S_{2}\wr S_{n}$. (Here, $\|\cdot\|$ denotes any norm on the space of $2n\times 2n$ matrices; for example, we may take $\|A\|=\max\{|a_{i,j}|:1\leq i,j\leq 2n\}$, the maximum absolute values of the entries of $A$.)
@@
Let $A\in\mathrm{Sp}(2n,\mathbb{Z})$ be such that the characteristic polynomial of $A$ is irreducible with Galois group $S_{2}\wr S_{n}$. If $n\geq 2$, then for every integer $m\geq 1$, the characteristic polynomial of $A^{m}$ is irreducible with Galois group $S_{2}\wr S_{n}$ and $A$ has a non-unit length eigenvalue. If $n=1$, this same conclusion holds as well, provided that the characteristic polynomial of $A$ is not $x^{2}+1$, $x^{2}+x+1$, or $x^{2}-x+1$.
@@
Suppose to the contrary that $\lambda^{m}=\alpha$ for some $\alpha\in F$. The extension $\mathbb{Q}(\lambda)/F$ is Galois (it is degree $2$), with the nontrivial automorphism sending $\lambda$ to $\frac{1}{\lambda}$. Applying this automorphism (which fixes $\alpha$) and simplifying, we also find that $\lambda^{m}=1/\alpha$. This implies that $\alpha=\pm 1$ and that $\lambda$ is a root of unity. This cannot be the case if $n\geq 2$, since the extension $\mathbb{Q}(\lambda)$ is not abelian (the group $S_{2}\wr S_{n}$ is not abelian for any $n\geq 2$), nor can it be the case if $n=1$ unless $f_{A}$ is one of the listed polynomials (the listed polynomials are the only irreducible quadratics whose roots are roots of unity).
Finally, we note that if an algebraic integer is not a root of unity, it must have a Galois conjugate off the unit circle. As $A$ cannot have eigenvalues that are roots of unity, it must have a non-unit length eigenvalue. This completes the proof. ∎
To prove Theorem B.1 in the case $m=1$, we will employ a strategy of Davis, Duke, and Sun [DDS], incorporating results of Chavdarov [Chavdarov] and Gorodnik and Nevo [GorodnikNevo]. Namely, this involves extending statistics for reciprocal polynomials to the case of symplectic matrices via equidistribution results in ergodic theory.
We begin with the following lemma from [DDS].
@@
[DDS]*Lemma 2 Let $n\geq 2$ and let $G$ be a subgroup of $S_{2}\wr S_{n}$ containing a $2$-cycle, 44-cycle, (2n2)(2n-2)-cycle, and a $2n$-cycle. Then $G=S_{2}\wr S_{n}$.
Lemma B.3 suggests that sieving is a natural approach to Theorem B.1. Specifically, the characteristic polynomial $f_{A}(x)$ of any element $A\in\mathrm{Sp}(2n,\mathbb{Z})$ must be reciprocal (i.e., palindromic), that is, it satisfies $f_{A}(x)=x^{2n}f_{A}(\frac{1}{x})$; let $\mathcal{P}_{2n}^{\mathrm{rec}}(\mathbb{Z})$ be the set of monic reciprocal polynomials in $\mathbb{Z}[x]$ of degree $2n$, so that $f_{A}\in\mathcal{P}_{2n}^{\mathrm{rec}}(\mathbb{Z})$. Moreover, the factorization type of $f_{A}\,(\mathrm{mod}\,\ell)$ for primes $\ell$ not dividing the discriminant of $f_{A}$ corresponds to the cycle type of an element in $\mathrm{Gal}(f_{A})$ (namely, the Frobenius element at $\ell$). Thus, for any prime $\ell$, let $\mathcal{P}_{2n}^{\mathrm{rec}}(\mathbb{F}_{\ell})$ be the set of monic reciprocal polynomials of degree $2n$ over $\mathbb{F}_{\ell}$, and, motivated by Lemma B.3, for any $1\leq k\leq n$ we let $\mathcal{P}_{2n}^{\mathrm{rec}}(F_{\ell};(2k))$ be the subset of $\mathcal{P}_{2n}^{\mathrm{rec}}(\mathbb{F}_{\ell})$ consisting of polynomials factoring as an irreducible polynomial of degree $2k$ times a product of 2n2k2n-2k distinct linear polynomials. Observe that $\mathcal{P}_{2n}^{\mathrm{rec}}(\mathbb{F}_{\ell})=\ell^{n}$.
@@
$\mathcal{P}_{2n}^{\mathrm{rec}}(\mathbb{F}_{\ell};(2k))=\frac{1}{2^{n-k+1}\cdot k\cdot(n-k)!}\ell^{n}+O(\ell^{n-1}).$
@@
In particular, Lemma B.5 shows that the characteristic polynomials of elements of $\mathrm{Sp}(2n,\mathbb{F}_{\ell})$ are approximately equidistributed in $\mathcal{P}^{\mathrm{rec}}_{2n}(\mathbb{F}_{\ell})$. (Recall that $|\mathrm{Sp}(2n,\mathbb{F}_{\ell})|=\ell^{2n^{2}+n}+\linebreak O(\ell^{2n^{2}+n-1})$ and $\mathcal{P}_{2n}^{\mathrm{rec}}(\mathbb{F}_{\ell})=\ell^{n}$.) However, to obtain Theorem B.1, it is necessary to know that elements of $\mathrm{Sp}(2n,\mathbb{Z})$ are essentially equidistributed when reduced $(\mathrm{mod}\,\ell)$, and that these reductions are essentially independent for different primes $\ell$. In other words, we are asking for an effective form of approximation for the group $\mathrm{Sp}(2n,\mathbb{Z})$. For this, we have the following consequence of work of Gorodnik and Nevo [GorodnikNevo].
@@
$\#\{A\in\mathrm{Sp}(2n,\mathbb{Z}):\|A\|\leq H,A\equiv A_{0}\,(\mathrm{mod}\,q)\}$
@@
[…]
@@
(B.1) $\{A\in\mathrm{Sp}(2n,\mathbb{Z}):\|A\|\leq H,f_{A}\in\mathcal{E}_{2k}\}=o(\#\{A\in\mathrm{Sp}(2n,\mathbb{Z}):\|A\|\leq H\}).$
@@
$\#\{A\in\mathrm{Sp}(2n,\mathbb{Z}):\|A\|\leq H,f_{A}\in\mathcal{E}_{2k}(q)\}$
@@
Acknowledgements. EK is supported by MathWorks and by NSF GRFP under grant No. 1745302. She would like to thank Semyon Dyatlov, who is partially supported by NSF CAREER grants DMS-1749858 and DMS-2400090, for suggesting and advising this project. She would also like to thank Bjorn Poonen and Vijay Srinivasan for their help with Proposition A.3. She is grateful to RJLO for his help with the Galois group explanation after Proposition A.3. She thanks Zeev Rudnick for his helpful comments on the first version of this paper. TCA is supported by NSF DMS-2231990 and NSF CAREER DMS-2237937. She thanks Peter Sarnak for introducing a variant of this problem to her. RJLO is supported by NSF DMS-2200760 and a Simons Fellowship in Mathematics. We are grateful to the anonymous referee for their numerous helpful comments on a previous version of this paper.